# Value function interference and greedy action selection in value-based multi-objective reinforcement learning

### 09 Feb 2024 v1

@@ Abstract
Multi-objective reinforcement learning (MORL) algorithms extend conventional reinforcement learning (RL) to the more general case of problems with multiple, conflicting objectives, represented by vector-valued rewards. Widely-used scalar RL methods such as Q-learning can be modified to handle multiple objectives by (1) learning vector-valued value functions, and (2) performing action selection using a scalarisation or ordering operator which reflects the user’s utility with respect to the different objectives. However, as we demonstrate here, if the user’s utility function maps widely varying vector-values to similar levels of utility, this can lead to interference in the value-function learned by the agent, leading to convergence to sub-optimal policies. This will be most prevalent in stochastic environments when optimising for the Expected Scalarised Return criterion, but we present a simple example showing that interference can also arise in deterministic environments. We demonstrate empirically that avoiding the use of random tie-breaking when identifying greedy actions can ameliorate, but not fully overcome, the problems caused by value function interference.
Keywords multi-objective reinforcement learning  $\cdot$ Q-learning  $\cdot$ value-based reinforcement learning  $\cdot$ action selection
@@ 1 Introduction and Background
Multi-objective reinforcement learning (MORL) is a form of reinforcement learning (RL) designed for problems with multiple, conflicting objectives (Roijers et al., 2013). Whereas conventional, single-objective RL algorithms assume that the agent is provided with a scalar reward after each action, in MORL these rewards are vectors with a separate component for each objective. Prior research has shown that applying single-objective methods to multiobjective tasks may result in unsuitable, inferior or inflexible solutions (Roijers et al., 2013, 2015), and so in recent years there has been considerable interest in developing MORL methods. In many cases these MORL algorithms are extensions of existing RL methods, including value-based approaches such as Q-learning (Gábor et al., 1998; Natarajan and Tadepalli, 2005; Van Moffaert et al., 2013; Van Moffaert and Nowé, 2014; Ruiz-Montiel et al., 2017; Vamplew et al., 2017a).
@@
    When action selection is performed, a greedy action must be identified in a way which is consistent with the user’s overall utility Roijers et al. (2013); Zintgraf et al. (2015).
The action selection can in most cases be implemented via applying a scalarisation operator which maps the vector-values to scalars, and then regarding the action with the highest scalarised value as being the greedy selection. Depending on the user’s utility preferences this might be a simple linear weighting or a non-linear operator, such as the Chebyshev distance [^1] We note that when a non-linear operator is used, then both the action selection and the Q-values learned by the agent must be conditioned not just on the current state of the environment, but also on some measure of the rewards received so far (Vamplew et al., 2021). An exception is in the case where rewards are zero except at terminal states (Issabekov and Vamplew, 2012). As this issue is orthogonal to those considered in this paper, for clarity we will only consider this latter form of environment within this paper - however the observations made here are equally applicable to the more general case.. For some forms of utility (e.g. those based on lexicographic ordering, which often arise in the context of meeting constraints (Gábor et al., 1998) or in the problems where fairness is a consideration (Siddique et al., 2020)), it may not be possible to represent the utility directly in terms of scalarisation (Debreu, 1997). In such cases, the Q-learning agent can instead use an ordering operator to identify the greedy policy.
Algorithm 1 provides a generic form of multi-objective Q-learning. The action-selection function $U$ may be based either on explicit scalarisation of value-vectors or on an ordering operator, depending on the nature of the user’s utility function.
@@
1:input: learning rate $\alpha$, discounting term $\gamma$, eligibility trace decay term $\lambda$, number of objectives $n$, action-selection function $U$ and any associated parameters
2:for all states $s$, actions $a$ and objectives $o$ do
3:     initialise $Q_{o}(s,a)$
@@
9:     sums of prior rewards $P_{o}$ = 0, for all $o$ in 1..$n$
10:     observe initial state $s_{t}$
11:     $s_{t}$ = $(s_{t},P)$triangleright create augmented state
12:      select $a_{t}$ from an exploratory policy derived using $f(Q(s))$
13:     for each step of the episode do
@@
15:         $P=P+R_{t}$
16:         $s_{t+1}$ = $(s_{t+1},P)$triangleright create augmented state
17:         $V(s_{t+1})=Q(s_{t+1})+P$triangleright create value vector
18:         select $a^{*}$ from a greedy policy derived using $U(V(s_{t+1}))$
19:         select $a^{\prime}$ from an exploratory policy derived using $U(V(s_{t+1}))$
20:         $\delta=R_{t}+\gamma Q(s_{t+1},a^{*})-Q(s_{t},a_{t})$
21:         $e(s_{t},a_{t})$ = 1
22:         for each state $s$ and action $a$ do
23:              $Q(s,a)=Q(s,a)+\alpha\delta e(s,a)$
@@
25:                  $e(s,a)=\gamma\lambda e(s,a)$
26:              end if
27:         end for
28:         $s_{t}=s_{t+1},a_{t}=a^{\prime}$
@@
30:end for
One further consideration in MORL is the choice of optimisation criteria, as there are two distinct options compared with just a single possible criteria in conventional RL (Roijers et al., 2013). The first one is to maximise the Expected Scalarised Return (ESR). In this approach, the agent aims to maximise the expected value which is first scalarised by the utility function, as shown below (Eq 1) where $w$ is the parameter vector for utility function $f$, $r_{k}$ is the vector reward on time-step $k$, and $\gamma$ is the discounting factor
$V^{\pi}_{\bf w}(s)=E[f(\sum^{\infty}_{k=0}\gamma^{k}\mathbf{r}_{k},{\bf w})\ |\ \pi,s_{0}=s)$ (1)
@@
$V^{\pi}_{\bf w}(s)=f({\bf V}^{\pi}(s),{\bf w})=f(E[\sum^{\infty}_{k=0}\gamma^{k}\mathbf{r}_{k}\ |\ \pi,s_{0}=s],{\bf w})$ (2)
@@
Multiobjective extensions of Q-learning have been amongst the most widely used forms of MORL reported so far in the literature (Gábor et al., 1998; Natarajan and Tadepalli, 2005; Van Moffaert et al., 2013; Van Moffaert and Nowé, 2014; Ruiz-Montiel et al., 2017; Vamplew et al., 2017a). However, as we will demonstrate via example and empirically in this paper, when used in combination with a non-linear utility function an issue which we will label value function interference can arise, which may slow learning or cause the agent to converge to a sub-optimal policy.
2 Interference in multi-objective value functions for deterministic environments
@@
We start by considering the simplest case, which are environments where the choice of starting state, the state transitions and the rewards are fully deterministic. Assuming that the policy being executed is also deterministic (as would usually be the case for methods based on extensions of Q-learning), then in this context the SER and ESR optimisation criteria are equivalent (every execution of a policy results in the same return, which of course is also the same as the mean return considered by SER).
2.1 An example of value function interference in a stochastic MOMDP
@@
As an illustrative example, consider the simple deterministic multi-objective Markov Decision Process (MOMDP) shown in Figure 1. This MOMDP consists of three states ($A$, $B$, and $C$). In each state two actions are available ($a_{1}$ and $a_{2}$). The MOMDP has three objectives, so the reward vector has three elements. There are four possible policies, as summarised in Table 1.
@@
Policy label Action in State A Action in State B Action in State C Vector return Utility
0 $a_{1}$ $a_{1}$ (7, -1, -5) 9
1 $a_{1}$ $a_{2}$ (7, -5, -1) 9
2 $a_{2}$ $a_{1}$ (8, -3, -3) 7
3 $a_{2}$ $a_{2}$ (0, -5, -5) -25
Table 1: The four policies for the MOMDP from Figure 1. The scalar utility in the final column assumes that the utility equals twice the first objective minus the product of the second and third objectives.
There are many ways in which a utility function could be derived from the rewards [^2] The actual specification of rewards and utilities is outside the scope of this project; we focus here simply on examining whether MORL methods will optimise for the utilities and rewards as defined by the user, irrespective of the structure of those definitions. The issues identified here could arise for any non-linear definition of utility., but let us assume that $U$ equals twice the value of the first objective minus the product of the second and third objectives (which are always negative). Under these settings $U(B,a_{1})=U(B,a_{2})=9,U(C,a_{1})=7$, and $U(C,a_{2})=-25$. Therefore clearly the optimal policy should be to select action $a_{1}$ in state $A$, followed by either action in state $B$ i.e. either policy 0 or policy 1 (this is true for both the SER and ESR criterion).
However the existence of two equally optimal actions in state $B$ creates an issue in the context of multi-objective rewards. The usual approach taken in Q-learning where more than one optimal action exists is to simply randomly select between them. However this random selection introduces an element of stochasticity into the vector rewards which the agent receives, which in turn leads to interference in the values which the agent learns. This interference problem does not arise in single-objective RL, because two actions which are equally optimal must, by definition, share the same mean scalar reward, whereas in the multi-objective case two actions may share the same utility while having differing mean vector rewards.

### 22 Apr 2026 v2

# Issues with Value-Based Multi-objective Reinforcement Learning: Value Function Interference and Overestimation Sensitivity

@@ Abstract
Multi-objective reinforcement learning (MORL) algorithms extend conventional reinforcement learning (RL) to the more general case of problems with multiple, conflicting objectives, represented by vector-valued rewards. Widely-used scalar RL methods such as Q-learning can be modified to handle multiple objectives by (1) learning vector-valued value functions, and (2) performing action selection using a scalarisation or ordering operator which reflects the user’s preferences with respect to the different objectives. This paper investigates two previously unreported issues which can hinder the performance of value-based MORL algorithms when applied in conjunction with a non-linear utility function – value function interference, and sensitivity to overestimation. We illustrate the nature of these phenomena on simple multi-objective MDPs using a tabular implementation of multiobjective Q-learning.
@@ 1 Introduction and Background
Multi-objective reinforcement learning (MORL) is a form of reinforcement learning (RL) designed for problems with multiple, conflicting objectives (Roijers et al., 2013; Hayes et al., 2022). Whereas conventional, single-objective RL algorithms assume that the agent is provided with a scalar reward after each action, in MORL these rewards are vectors with a separate component for each objective. Prior research has shown that applying single-objective methods to multi-objective tasks may result in unsuitable, inferior or inflexible solutions (Roijers et al., 2013; 2015), and so in recent years there has been considerable interest in developing MORL methods. In many cases these MORL algorithms are extensions of existing RL methods, including value-based approaches such as Q-learning (Gábor et al., 1998; Natarajan and Tadepalli, 2005; Van Moffaert et al., 2013; Van Moffaert and Nowé, 2014; Mossalam et al., 2016; Vamplew et al., 2017b; Abels et al., 2019).
@@
    When action selection is performed, a greedy action must be identified in a way which is consistent with the user’s overall utility (Roijers et al., 2013; Zintgraf et al., 2015).
The action selection can in most cases be implemented via applying a scalarisation operator which maps the vector-values to scalars, and then regarding the action with the highest scalarised value as being the greedy selection. Depending on the user’s utility preferences this might be a simple linear weighting or a non-linear operator, such as the Chebyshev distance. For some forms of utility (e.g. those based on lexicographic ordering, which often arise in the context of meeting constraints (Gábor et al., 1998) or in problems where fairness is a consideration (Siddique et al., 2020)), it may not be possible to represent the utility directly in terms of scalarisation (Debreu, 1997). In such cases, the Q-learning agent can instead use an ordering operator to identify the greedy action. For simplicity in this paper we will restrict our examples to utilities which can be implemented via scalarisation.
Algorithm 1 provides a generic form of multi-objective Q-learning. The algorithm presented here is similar to that previously described in Vamplew et al. (2021b) and Ding et al. (2025), but has been modified to correctly handle discounting of future rewards. The agent maintains a vector $P$ of the discounted return accumulated so far (Line 14). This is combined with appropriately discounted Q-values for the current state to estimate the complete discounted return $V$ (Line 16), which is subsequently used in conjunction with $U$ to perform action selection (Lines 17-18).
Note that when a non-linear $U$ is used, then both the action selection and the Q-values learned by the agent must be conditioned not just on the current state of the environment, but also on some measure of the rewards received so far (Vamplew et al., 2021b). [^1] An exception is in the case where rewards are zero except at terminal states. (Issabekov and Vamplew, 2012). For clarity the environments used in this paper do have this simplifying property. However the observations made here are equally applicable to the more general case. This is the purpose of the augmented state created on Lines 10 and 15. Again, the algorithm specified here introduces a modification from that reported in the previous literature, in that the augmented state concatenates the environment state information with both the accumulated return vector $P$, and also the time-step. The need for the latter term has not to our knowledge previously been identified in the literature on multi-objective Q-learning – while Cai et al. (2023) do include $t$ within their definition of state (see footnote 3 on page 3 of their paper), they state that this is required only for the finite horizon setting. However it can be seen from Algorithm 1 that the value of $V$, and therefore the choice of action, depends on $t$ as this is used to determine the amount of discounting applied to the Q-values. It is possible that the structure of a MOMDP may allow for the agent to reach a particular environmental state $s$ with the same accumulated return $P$ after different numbers of time-steps. In this case the optimal policy may differ depending on the time-step, and hence this must be included in the augmented state $S^{A}$. Without this the policy may be non-stationary with respect to $S^{A}$ and the agent may fail to converge to a stable policy. In practice as the value of $t$ is potentially unbounded, it may be preferable to use $\gamma^{t}$ in the augmented state as this will naturally be normalised in the range $0\ldots 1$].
@@
1:input: learning rate $\alpha$, discounting term $\gamma$, eligibility trace decay term $\lambda$, number of objectives $n$, scalarisation function $U$ and any associated parameters
2:for all augmented states $s^{A}$, actions $a$ and objectives $o$ do
3:  initialise $Q_{o}(s^{A},a)$
@@
9:  initialise sum of prior rewards $P$ to a zero vector
10:  observe initial state $s_{0}$
11:  $s^{A}_{0}=(s_{0},P,\gamma^{0})$ $\triangleright$ create augmented state
12:   select $a_{0}$ from an exploratory policy derived using $U(Q(s^{A}_{0}))$
13:  for each step $t$ of the episode do
@@
15:   $P=P+\gamma^{t}R_{t}$
16:   $s^{A}_{t+1}=(s_{t+1},P,\gamma^{t+1})$ $\triangleright$ create augmented state
17:   $V(s^{A}_{t+1})=P+\gamma^{t+1}Q(s^{A}_{t+1})$ $\triangleright$ create value vector for all actions
18:   select $a^{*}$ from a greedy policy derived using $U(V(s^{A}_{t+1}))$
19:   select $a^{\prime}$ from an exploratory policy derived using $U(V(s^{A}_{t+1}))$
20:   $\delta=R_{t}+\gamma Q(s^{A}_{t+1},a^{*})-Q(s^{A}_{t},a_{t})$
21:   $e(s^{A}_{t},a_{t})=1$
22:   for each augmented state $s^{A}$ and action $a$ do
23:     $Q(s^{A},a)=Q(s^{A},a)+\alpha\delta e(s^{A},a)$
@@
25:      $e(s^{A},a)=\gamma\lambda e(s^{A},a)$
26:     else
27:      $e(s^{A},a)=0$
28:     end if
@@
30:   $a_{t+1}=a^{\prime}$
31:  end for
32:end for
One further consideration in MORL is the choice of optimisation criteria, as there are two distinct options compared with just a single possible criteria in conventional RL (Roijers et al., 2013). The first one is to maximise the Expected Scalarised Return (ESR). In this approach, the agent aims to maximise the expected value which is first scalarised by the utility function, as shown below (Eq 1) where $w$ is the parameter vector for utility function $U$, $r_{k}$ is the vector reward on time-step $k$, and $\gamma$ is the discounting factor
$V^{\pi}_{\bf w}(s)=E[U(\sum^{\infty}_{k=0}\gamma^{k}\mathbf{r}_{k},{\bf w})\ |\ \pi,s_{0}=s]$ (1)
@@
$V^{\pi}_{\bf w}(s)=U({\bf V}^{\pi}(s),{\bf w})=U(E[\sum^{\infty}_{k=0}\gamma^{k}\mathbf{r}_{k}\ |\ \pi,s_{0}=s],{\bf w})$ (2)
@@
One further distinction between MORL algorithms is whether they are single-policy or multi-policy in nature (Vamplew et al., 2011; Roijers et al., 2013). The former aims to learn a single policy which is optimal for a specific choice of utility parameters, whereas the latter learns a set of policies where each policy is optimal for a different choice of parameters for the utility function. For clarity this paper restricts its focus to single-policy algorithms, but the issues identified here would also arise in value-based multi-policy algorithms.
@@
Multi-objective extensions of Q-learning have been amongst the most widely used forms of MORL reported so far in the literature (Gábor et al., 1998; Natarajan and Tadepalli, 2005; Van Moffaert et al., 2013; Van Moffaert and Nowé, 2014; Ruiz-Montiel et al., 2017; Vamplew et al., 2017b). However, as we will demonstrate via example and empirically in this paper, when used in combination with a non-linear utility function two issues can arise which may significantly slow learning or cause the agent to converge to a sub-optimal policy. Section 2 examines an issue which we label value function interference where the expected vector-values learned by the agent may lead to sub-optimal action selection. We show that this problem can occur in both stochastic and deterministic environments, and examine a modified form of action-selection which can mitigate (though not eliminate) the problem in the latter case. Section 3 examines the impact of overestimation of Q-values on MORL. It is well known from prior studies of scalar RL that methods such as Q-learning can be prone to overestimation (Thrun and Schwartz, 1993; Van Hasselt et al., 2016). Here we identify for the first time that this issue may be even more impactful on learning in the context of multiple objectives.
2 Interference in multi-objective value functions
As discussed in the Introduction, extensions of value-based RL methods such as Q-learning to MORL require learning vectors for the Q-values. In this type of algorithm, the vector-values learned for a particular state-action pair represent the expected return averaged over multiple different future returns (due to stochasticity in either the environment and/or the policy of the agent). When used in combination with a non-linear utility function $U$, the utility of this expected return vector may differ from that of the returns from which it is derived, and this may lead to action-selection which is not optimal with regards to $U$. We label this phenomenon value function interference.
The graphs in Figure 1 illustrate a set of circumstances under which value function interference may arise and cause learning of a sub-optimal policy. For simplicity of visualisation, this figure illustrates the case of a single-objective reward. While the application of a utility function is not standard practice in single-objective RL, it is feasible and potentially beneficial to do so (Vamplew et al., 2024). In both graphs, we assume the existence of two actions $a_{1}$} and $a_{2}$. We assume that $a_{1}$ stochastically produces two outcomes, with returns denoted $Q(a_{1})^{low}$ and $Q(a_{1})^{high}$, and that these occur with equal probability such that $Q(a_{1})=(Q(a_{1})^{low}+Q(a_{1})^{high})/2$. Meanwhile $a_{2}$ is assumed to produce a deterministic outcome [^2] We note that value function interference may still arise if the outcomes of $a_{2}$ are stochastic, as long as the range of variation is lower than for $a_{1}$. denoted $Q(a_{2})$.
Figure 1: An illustration of value-function interference for both a convex and concave utility function. In both cases action $a_{1}$ has a stochastic return where $Q(a_{1})^{low}$ and $Q(a_{1})^{high}$ occur with equal probability. The highlighted region indicates conditions under which the agent may prefer the incorrect action with regards to ESR optimality.
As $a_{2}$ has deterministic outcomes, its utility will be the same under both the SER and ESR criterion, and equal to $U(Q(a_{2}))$. However as $a_{1}$ outcomes are stochastic, its utility will differ under these criterion. For SER, the utility is $U(Q(a_{1}))$ which equals $U(Q(a_{1})^{low}*0.5+Q(a_{1})^{high}*0.5)$ (i.e. the estimated per-objective returns are probabilistically combined over all outcomes, and then the utility function is applied. This is the value which would be learned and used within Algorithm 1. In contrast, for ESR the relevant utility is $U(Q(a_{1})^{low})*0.5+U(Q(a_{1})^{high})*0.5$ (i.e. a probabilistically weighted average of the scalarised return of each outcome). For non-linear definitions of $U$, these values will not be equal. Figure 1 shows that for a convex utility function (left) $SER(a_{1})<ESR(a_{1})$ whereas for a concave utility function (right) $ESR(a_{1})<SER(a_{1})$.
This issue may lead to selection of actions which are incompatible with optimising the ESR criterion, if another action ($a_{2}$) has an expected return falling within the shaded region in these graphs. In the case on the left, if $SER(a_{1})<U(Q(a_{2}))<ESR(a_{1})$ then $a_{2}$ will incorrectly be identified as the optimal action whereas $a_{1}$ is preferable under ESR. Conversely in the case on the right, if $ESR(a_{1})<U(Q(a_{2}))<SER(a_{1})$ then $a_{1}$ will incorrectly be identified as optimal action whereas $a_{2}$ is preferable under ESR.
2.1 Value function interference in stochastic MOMDPs
We start by considering environments with some element of stochasticity (in either the state transitions or rewards), as this is perhaps the more obvious context in which value function interference can occur. For stochastic environments the SER and ESR criteria are not equivalent, and value interference issues may arise for ESR-optimal learning which do not occur for SER-optimal learning.
Consider the very simple multi-objective Markov Decision Process (MOMDP) [^3] In fact, as this MOMDP has only a single non-terminal state, this example demonstrates that value function interference can arise even for ESR optimisation of multi-objective multi-armed bandits (MOMABs). shown in Figure 2. There are many ways in which a utility function could be derived from the rewards in this MOMDP, but let us assume that $U$ equals twice the value of the first objective minus the product of the second and third objectives (which are always negative) [^4] The actual specification of rewards and utilities is outside the scope of this project; we focus here simply on examining whether MORL methods will optimise for the utilities and rewards as defined by the user, irrespective of the structure of those definitions. The issues identified here could arise for any non-linear definition of utility.. Action $a_{1}$ has a stochastic reward which varies between (7, -1, -5) and (7, -5, -1) with equal probability. The utility of this action differs depending on whether we are optimising with respect to ESR or SER. $U_{SER}(a_{1})$=5 (the scalarised value of the expected vector reward) while $U_{ESR}(a_{1})$=9 (the expected scalarised value, which is 9 for both of the possible outcomes). Action $a_{2}$ deterministically returns a reward of (8, -3, -3) – as this is deterministic $U_{SER}(a_{2})=U_{ESR}(a_{2})$ = 7.
Therefore the SER-optimal policy in this case is to select $a_{2}$, while the ESR-optimal policy selects $a_{1}$. The SER-optimal policy can be determined on the basis of the vector Q-values learned by the agent, but this is not true for the ESR-optimal policy, as these values only represent the expected vector return which is insufficient for the agent to determine the per-episode expected scalarised return.
To demonstrate the impact of value function interference, experiments were run on this MOMAB using a tabular multi-objective Q-learning agent. A hyperparameter sweep was performed across different values for the learning rate ($\alpha$) and starting exploration parameter (using $\epsilon$-greedy exploration with $\epsilon$ linearly decayed to zero over the trial). The $\lambda$ and $\gamma$ hyperparameters were fixed at 0.95 and 1 respectively, and to produce more directed exploration Q-values were optimistically initialised to (12,0,0). For each set of hyperparameters 100 trials of each agent were executed. Each trial lasted for 500 episodes, and at the end of the trial the final greedy policy of the agent was identified. Figure 3 shows the final greedy arm selection over 100 trials of each hyperparameter setting. It clearly highlights the problems caused by value function interference, with all but one trial using a learning rate $\alpha\leq 0.7$ converging to the incorrect arm, thereby causing a large loss in utility. Only trials using a very high learning rate were successful, and such high learning rates are unlikely to work effectively for more complex stochastic MOMDPs.
Figure 3: Heatmaps showing the frequency with which the correct arm $a_{1}$ (left) and incorrect arm $a_{2}$ (right) were selected as the final greedy arm for the stochastic MOMAB across 100 trials using different hyperparameter settings.
2.2 Value function interference in deterministic MOMDPs
@@
Now consider the case of environments where the choice of starting state, the state transitions and the rewards are fully deterministic. Assuming that the policy being executed is also deterministic, then in this context the SER and ESR optimisation criteria are equivalent (every execution of a policy results in the same return, which of course is also the same as the mean return considered by SER). Given the fully deterministic characteristics of this setting, it is less clear how value function interference might arise. However we will demonstrate that it can still occur, due to two different underlying causes.
@@
Value function interference manifests differently in the deterministic case. Rather than being evident in the values learned for the state-action pair where stochastic rewards occur as in Section 2.1, in this context issues arise with the values learned for predecessor states. Therefore rather than a MOMAB, we will require a multi-state MOMDP to demonstrate this issue. As an illustrative example, consider the simple deterministic MOMDP shown in Figure 4. This MOMDP consists of three states ($A$, $B$, and $C$). In each state two actions are available ($a_{1}$ and $a_{2}$). The MOMDP has three objectives, so the reward vector has three elements. There are four possible deterministic stationary policies, as summarised in Table 1.
@@
Policy label Action in Action in Action in Vector return Utility
State A State B State C
0 $a_{1}$ $a_{1}$ - (7, -1, -5) 9
1 $a_{1}$ $a_{2}$ - (7, -5, -1) 9
2 $a_{2}$ - $a_{1}$ (8, -3, -3) 7
3 $a_{2}$ - $a_{2}$ (0, -5, -5) -25
Table 1: The four policies for the deterministic MOMDP from Figure 4. The scalar utility in the final column assumes that the utility equals twice the first objective minus the product of the second and third objectives.
Assuming the same utility function as used previously, $U(B,a_{1})=U(B,a_{2})=9,U(C,a_{1})=7$, and $U(C,a_{2})=-25$. Therefore clearly the optimal policy should be to select action $a_{1}$ in state $A$, followed by either action in state $B$ i.e. either policy 0 or policy 1 (this is true for both the SER and ESR criterion).
However the existence of two equally optimal actions in state $B$ creates an issue in the context of multi-objective rewards. The usual approach taken in Q-learning where more than one optimal action exists is to simply randomly select between them. However this random selection violates our earlier assumption that the policy is deterministic. This introduces an element of stochasticity into the vector rewards which the agent receives, which in turn leads to interference in the values which the agent learns. This interference problem does not arise in single-objective RL, because two actions which are equally optimal must, by definition, share the same mean scalar reward, whereas in the multi-objective case two actions may share the same utility while having differing mean vector rewards.